Reports 1-1 of 1 Clear search Modify search
takafumi.ushiba - 12:32 Tuesday 07 December 2021 (19107) Print this report
ETMX BF V2 LVDT signal became noisy after the work on 30 of November

I started to check the spectra of ITMX and ETMX for health check and found the ETMX BF LVDT V2 noise is significantly larger than the reference.
After some investigation, something happened on 30 of November because the noise on 30 of November was the same as the reference while the noise on 1 of December was larger.

I started to check the spectra of ITMX and ETMX sensors and found the problem on ETMX BF LVDT V2 signal.
Figure 1 shows the spectra of ETMX BF LVDTs and we can see significant difference on left midle spectra (BF LVDT V2).

Then, I checked from when the spectra was strange and found that it was after the work on November 30.
Figure 2 shows the spectra on 30 of November at 6 am, which had the same noise level as the reference (the spectra at low frequncies quite different because the suspension was clamped but we can still compare the noise level by looking at high frequency region).
Figure 3 shows the spectra on 1 od December at 6 am, which had significantly larger noise compared to the reference.

Since we had a lot of work on 30 of November at X end (we touched both IP and BF stage), we cannot confirm the place which is likely to have the problem from the working log.
So, I asked Takahashi-san to investigate in this afternoon.

Images attached to this report
Comments to this report:
ryutaro.takahashi - 21:20 Tuesday 07 December 2021 (19120) Print this report

[Takahashi, Sato]

We investigated the noise in the BF V2 LVDT. In the work on 30 Nov., we moved the secondary coils. So the cable connection around the secondary coils might be changed. The resistances of the LVDT coils were the same as the measurement on 25 Nov. within 1 ohm, and the resistances were not changed when we shook the cables to the coils. The V2 spectrum itself is not so different from the V2 and V3 spectra. The increased noise may be due to the present DC level of V2 output  (12600 count) being much larger than the previous DC level (-1400 count).

Images attached to this comment
Search Help